Sap Paper Mill is considering $165 million upgrade of its machinery. Once the plant is upgraded, the machinery will last for 30 years. The upgrade is expected to generate the following net cash flows: $25 million annually in the first decade after the investment; $20 million annually in the second decade after the investment, and $15 million annually in the last decade of the investment. At the end of the 30th year, the firm also anticipates that it can sell the equipment (as parts) for a total salvage value of $1 million.   



Before committing to this investment, the company asked you, their Business Consultant, to evaluate this capital budgeting project given the various sources of funding available to them. To assist you, the firm gathered the following data:



The firm's corporate tax rate is currently 40 percent. The tax rate is not expected to change throughout the life of the project.
The company can issue 30-year bonds with a 12% coupon, paid semi-annually, and par value of $1,000 for $1,153.72. The bond’s flotation costs would be 1% of the par value.
The current price of the firm's 10 percent, $100 par value, preferred stock is $113.10. If the firm issues new preferred stock, the firm will have to maintain the preferred stock dividend rate constant, and pay $2 in flotation costs per each new preferred share issued.
The firm’s common stock is currently selling for $50 per share. Its last dividend (D0) was $4.19 per share. Common stock dividends are expected to grow at a constant rate of 5 percent per year in the foreseeable future.
If the firm issues new common stock, the net proceeds per share are estimated to be 15 percent less than the current price of common share.
Sap’s beta is 1.2. The yield on a 30-year US Treasury Bond is 7 percent per year, and the S&P’s rate of return is currently 13 percent per year.
The firm’s target capital structure is 30 percent long-term debt, 10 percent preferred stock, and 60 percent common stock equity.


Given this information, the firm asked you to answer the following questions and provide supporting evidence (formulas, calculations, etc) behind your evaluations.

What cost(s) of capital should the firm use when evaluating whether or not to invest in this project? Why should they use this/these cost(s) of capital? [10 points]
They should use the weighted average cost of capital because
What is the firm’s after-tax cost of debt financing?
What is the firm’s cost of preferred stock financing? [10 points]
According to CAPM, what is the firm’s cost of financing using the retained earning? [10 points]
According to the Gordon Growth Model, what is the firm’s cost of financing using the retained earning? [10 points]
What is the firm’s cost of equity, if the firm issues new common stock? [10 points]
What is the firm’s weighted average cost of capital if the firm uses debt, preferred stock and retained earnings? What cost of retained earnings did you use here and why?
What is the firm’s weighted average cost of capital if the firm uses debt, preferred stock and equity from the new common stock issue? [15 points]
Should the firm finance the plant’s machinery upgrade? If so, how should the firm finance it? Provide evidence to support your recommendation [10 points]
Answer #1

1.They should use the weighted average cost of capital because

Given that "the firm’s target capital structure is 30% long-term debt, 10 % preferred stock, and 60% percent common stock equity"--it will be only proper if the costs of all types of capital , in proportion to their weights, are included --so as to assess the feasibility of the project , in light of the costs of financing it.

2.Firm’s after-tax cost of debt financing

We equate the net proceeds of the issue,ie. After flotation cost--to the PV s of the bond's cash flows, at the yield ,which is the before-tax cost of the bond.
ie. Net proceeds=(Pmt.*(1-(1+r)^-n)/r)+(FV/(1+r)^n)
where net proceeds=Issue price-Flotation cost,ie. 1153.72-(1%*1000)=1143.72
Pmt.= the periodic coupon pmt., ie. 1000*12%/2= 60
r= the semi-annual yield --to be found out --??
n= no.of semi-annual coupon periods=30*2= 60
FV=Face value= $ 1000
Now, plugging in the values in the formula,
1143.72=(60*(1-(1+r)^-60)/r)+(1000/(1+r)^60)
& solving for r, we get the semi-annual before tax cost as,
5.21%
So. Annual before-tax cost=(1+5.21%)^2-1=
10.69%
After-tax cost of the bond= Before-tax cost*(1-Tax Rate)
ie. 10.69%*(1-40%)=
6.41%

3.Firm’s cost of preferred stock financing

Cost of preferred stock=$ dividend/stock/Net proceeds per stock
Net proceeds /stock=Issue price-Flotation cost
(10%*100)/(113.1-2)=
9%

4.According to CAPM,Firm’s cost of financing using the retained earning

As per CAPM,Cost of ret. Earn.k re=RFR+(Beta*(mkt.return-RFR))
ie.k re=7%+(1.2*(13%-7%))
14.2%

5.According to the Gordon Growth Model, Firm’s cost of financing using the retained earning

According to GGM, K re= (Next dividend/Current stock price)+Growth Rate
Where next dividend=(D0*(1+g))
k re=(4.19*(1.05)/50)+5%=
13.80%

6.Firm’s cost of equity, if the firm issues new common stock

we need to consider the flotation cost on new issue, & net the proceeds.
in which case, we can only use the DDM
ie.ke=(4.19*(1.05)/(50*(1-15%)))+5%=
15.35%

7.Firm’s weighted average cost of capital if the firm uses debt, preferred stock and retained earnings.

Cost of retained earnings to be used here is as per CAPM, as there is no new issue ,that is dependent on dividend cash flows.
So, the WACC=(Wt.d*kd)+(Wt. ps*k ps)+(Wt. re*k re)
ie.(30%*6.41%)+(10%*9%)+(60%*14.2%)=
11.34%

8.Firm’s weighted average cost of capital if the firm uses debt, preferred stock and equity from the new common stock issue

WACC=(Wt.d*kd)+(Wt. ps*k ps)+(Wt.e*k e)
ie.(30%*6.41%)+(10%*9%)+(60%*15.35%)=
12.03%
9. NPV of the upgrade at the WACC (11.34%) with new equity=
-165+(25*5.80620)+(20*1.98326)+(15*0.67744)+(1*0.03985)=
30.02
Millions
10. NPV of the upgrade at the WACC(12.03%) with Retained Earnings=
-165+(25*5.64329)+(20*1.81213)+(15*0.58190)+(1*0.03311)=
21.09
Millions.
11. YES. The firm can upgrade the plant as NPVs of upgrading at both WACC are POSITIVE & bring value to the firm.
NOTE: PV Factors used:
P/A,i=11.34%n=1-10 yrs. 5.8062
11-20 yrs. 1.98326
20-30 yrs. 0.67744
P/F yr.30 0.03985
P/A,i=12.03%n=1-10 yrs. 5.64329
11-20 yrs. 1.81213
20-30 yrs. 0.58190
P/F yr.30 0.03311